# Transparent AI Disclosure Obligations: Who, What, When, Where, Why, How

### 11 Mar 2024 v1

@@ Abstract.
EU AI Act, Article 52, generative artificial intelligence, disclosures, transparency, questions
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Human-centered computing Human computer interaction (HCI)
@@ 1. Introduction
“Fake realities will create fake humans. Or, fake humans will generate fake realities
and then sell them to other humans, turning them, eventually, into forgeries of themselves.”
@@
The field of Artificial Intelligence (AI), and the media sector at large, are undergoing a transformative change. Foundation models, developed on the basis of deep neural networks and self-supervised learning, have gained widespread acceptance (Bommasani et al., 2021). These models have led to the emergence of Generative AI (GenAI) tools like Midjourney and ChatGPT, and have shown impressive proficiency in producing media content, including images (Saharia et al., 2022), text (OpenAI, 2023), videos (Ho et al., 2022), and audio (Borsos et al., 2022). These have the potential to drastically impact users and the media sector, essentially blurring the line between fiction and reality as users engage with media. Model output has reached a level by which humans can no longer perceive GenAI outputs as distinguishable from human-generated content (Groh et al., 2023). What digital media appears to be true and authentic cannot necessarily be trusted, and is reportedly fueling the spread of mis- and disinformation [^1] https://www.axios.com/2023/02/21/chatbots-misinformation-nightmare-chatgpt-ai. Therefore it is imperative to disclose the use of AI in the generation/manipulation of media content. According to the World Economic Forum’s 2024 Global Risks Report (2024, 2024), “misinformation and disinformation” were ranked as the highest global risk anticipated over the next two years. This also comes at a time where OpenAI issues a statement regarding the upcoming 2024 worldwide elections (OpenAI, 2024), emphasizing: “Transparency around AI-generated content”. In fact, in the European Union (EU), such an obligation will be incorporated into the upcoming AI Act (Act, 2022). Despite that the AI Act is still in development and subject to change, the core obligation to “disclose AI-driven interaction” appears to be a stable one.
@@
Algorithmic approaches to automatically detect text (co-)produced using GenAI tools (e.g., ChatGPT) are so far unreliable (Sadasivan et al., 2023), and can result in consistent biases against specific user groups (e.g., non-native English writers (Liang et al., 2023)). While measures to encode provenance cryptographically in images (cf., Coalition for Content Provenance and Authenticity (C2PA, 2024)) or audio (SynthID (DeepMind, 2024)) are being taken, these are not widely implemented. Moreover, it remains unclear how these need to be displayed to users. Despite concerted efforts worldwide, constructing even simple disclosure measures (e.g., watermarks) that actually succeed in reducing public risks is itself a difficult task. Disclosure is defined as “to make known or public” (Webster, 2024), where in the context of AI, disclosures can “contain information about the data collection, data processing, and decision-making practices of a digital product and are voluntarily provided by the product’s vendor (an individual developer or an organization).” (Open Ethics Initiative (Lukianets, [n.d.])). A fundamental challenge here is in ensuring effective transparency measures, given the rapid pace of AI development and deployment. To mitigate harms and risks, regulatory efforts such as the AI Disclosure Act of 2023 in the United States (Torres, 2023) and the EU AI Act (Act, 2022) (currently under discussion (of the EU, [n.d.])) aim to tackle this. In this work, we focus on the EU AI Act proposition (Article 52: “Transparency obligations for providers and users of certain AI systems”), that addresses the issue of AI system transparency [^2] Even though the final text of the AI Act must still be agreed upon, the inclusion of a ”disclosure obligation” is (almost) a certitude..
@@ 2.1. EU AI Act’s Article 52 and AI disclosure implications
We focus on the EU AI Act (Act, 2022) (currently under discussion (of the EU, [n.d.])) proposition (Article 52: “Transparency obligations for providers and users of certain AI systems”), that addresses AI system transparency. It raises two important clauses:
@@
From these clauses, the types of disclosure are unspecified, and terms such as ‘authentic’ are undefined, leaving them open for interpretation. For example, ‘what’, ‘how’, or ‘when’ should such AI disclosures take place, even if automatically detected, remains unclear. While the most recent draft of the AI Act (Monitor, [n.d.]) touches on aspects of this, where Art. 52(3) (second subparagraph) highlights when to disclose, much remains uncharted regarding human oversight processes when exceptions are made. We believe it is important to carefully assess the AI Act’s disclosure implications, before proceeding to Human-AI interactions, wherein explainability (Panigutti et al., 2023) may be an additional challenge. Furthermore, authenticity is itself a multidimensional concept (Kernis and Goldman, 2006), which raises the question of what aspects of inauthentic media need to be disclosed. AI disclosures, be they automatic (algorithmic) or provided by system designers or other stakeholders, may include the following: revealing when content is AI-generated, registering these emerging AI systems with a database, summarizing copyrighted material used in training these systems, publication of risk assessments, or even trust certification labels (Scharowski et al., 2023). Importantly, encounters with AI-generated content can impact the human experience of algorithms, and more broadly the psychology of Human-AI interaction (Sundar, 2020), to which we turn to next.
@@ 3.2. Workshop materials
The workshop setup consisted of six flip charts that were attached to the walls of the workshop area (see Figure 1(b)). Each flip chart was given a heading of one of the 5W1H terms to indicate to participants which type of questions to ask there. Post-its and markers were provided. Article 52 and Article 3 were printed on an A4 landscape paper, and placed near each flip chart for quick reference. Article 3 provides definitions of ‘provider’ and ‘user’ [^3] In the latest draft version of the EU AI Act (leaked online on January 21, 2024) (Monitor, [n.d.]), ‘user’ is now referred to as ‘deployer’. However, this does not impact our workshop outcomes, even if participants interpreted ‘user’ as intended by the law (‘deployer‘) or as ‘recipient’., that helped participants have a clear understanding of what these ‘legal’ terms mean. Article 3 (current revision; bold added for emphasis) states:
@@ 3.3. Study procedure
Sessions began with a short  10 minute introduction to set the workshop context, to obtain participant consent (data protection and privacy), and for them to fill in a demographics form. Each workshop session consisted of three tasks, each lasting around 25-30 min., amounting to approximately one and a half hours in total. In the introduction, we first presented the EU AI Act, along with Article 52 and a brief explanation of disclosures. Participants were assigned an ID to note down with each post-it note. Participants were split into groups of two or three, depending on the total size in the session. The grouping was only for discussion and logistical purposes, so each question posed was still done individually. Task 1 and 2 were question generation tasks, where participants were instructed to individually generate questions for each of the 5W1H questions. After 10 minutes, they moved to a different question, such that at the end of  60 minutes each group has had the chance to provide their input for all six question types. Participants were also instructed to review the questions that were placed by participants on the flip chart, and to optionally vote with a +/- 1 post-it on the question if they were in agreement/disagreement. For Task 3, we divided each flip chart into two parts: users and providers. This provided participants the opportunity to reflect on and consider whether the posted questions were relevant for the user, the provider, or both (in which case it was placed in the center of the flip chart). Participants were offered candy in the session. At the end of each session, participants were thanked for their time and efforts.
@@ 4.4. Theme 4: Trust, Authenticity, and User Empowerment
A key theme (N=33) of our workshops was on trust, authenticity, and how users can be empowered. This consisted of four sub-themes. The first sub-theme includes questions on Authenticity, Provenance, and Transparency: “How important is it for users to know the authenticity of media?[Q81]; “How are we going to verify authenticity?[Q82]; “Where can people manually verify if it is AI generated?[Q84]. The second sub-theme tackles the Human-centered AI issue of Building Trust in AI Systems: “How will different users interpret signals differently?[Q89]; “How does disclosure impact people’s trust and perceived credibility?[Q91]; “How to responsibly disclose artistic uses of AI?[Q92]. The third sub-theme focuses on Empowering Users through Education & Awareness: “How can we train users to spot authentic content without relying on explicit cues?[Q99]; “What type of information empowers/triggers “act of resistance”?[Q100]; “Where should we get ask if we suspect things (images, texts) are generated by AI?[Q103]. Finally, the last sub-theme focuses on User Responsibility and Agency: “Why do users need to know they are interacting with AI?[Q110]; “What information should the users acquire? Are they responsible for it?[Q111]; “Why is it important for promoting democracy?[Q112]. The sub-themes here touch upon important prior work that aims to address authenticity and trust, in the context of media. Scharowski et al. (Scharowski et al., 2023) examined the potential for AI certification labels (e.g., the ”Digital Trust Label” by the 2023 Swiss Digital Initiative), and found that these can mitigate data-related concerns expressed by end-users (e.g., privacy and data protection), however other concerns (e.g., model performance) remain challenging to address. Longoni et al. (Longoni et al., 2022) found that people rated news headlines written by AI as less accurate than those written by humans. Recently, Toff & Simon (Toff and Simon, 2023) found that on average audiences perceive news labeled as AI-generated as less trustworthy, even when the articles themselves are not evaluated as any less accurate or unfair. However, these effects largely stem from those whose pre-existing levels of trust in news are higher to begin with and among those knowledgeable about journalism.
@@ 5.3. Conclusion
Our work contributes 149 questions clustered into five themes and 18 sub-themes, that we believe can assist us in tackling the challenge of transparently communicating AI disclosures – for media providers, as well as end-users consuming media content everyday across platforms and devices. We hope our questions underscore the importance for better understanding user needs and reactions to transparency obligations, and support establishing user-centric designs for transparent AI disclosures that ultimately foster democratic societies based on truth rather than AI-generated fiction.

### 13 Mar 2024 v2

@@ Abstract.
EU AI Act, Article 52, generative artificial intelligence, disclosures, transparency, obligations, law, research questions
journalyear: 2024copyright: rightsretainedconference: Extended Abstracts of the CHI Conference on Human Factors in Computing Systems; May 11–16, 2024; Honolulu, HI, USAbooktitle: Extended Abstracts of the CHI Conference on Human Factors in Computing Systems (CHI EA ’24), May 11–16, 2024, Honolulu, HI, USAdoi: 10.1145/3613905.3650750isbn: 979-8-4007-0331-7/24/05ccs: Human-centered computing Human computer interaction (HCI)
@@ 1. Introduction
“Fake realities will create fake humans. Or, fake humans will generate fake realities and then sell them to other humans, turning them, eventually, into forgeries of themselves.”
@@
The field of Artificial Intelligence (AI), and the media sector at large, are undergoing a transformative change. Foundation models, developed on the basis of deep neural networks and self-supervised learning, have gained widespread acceptance (Bommasani et al., 2021). These models have led to the emergence of Generative AI (GenAI) tools like Midjourney and ChatGPT, and have shown impressive capabilities in producing media content, including images (Saharia et al., 2022), text (OpenAI, 2023), videos (Ho et al., 2022), and audio (Borsos et al., 2022). These have the potential to drastically impact users and the media sector, essentially blurring the line between fiction and reality as users engage with media. Model output has reached a level by which humans can no longer perceive GenAI outputs as distinguishable from human-generated content (Groh et al., 2023). What digital media appears to be true and authentic cannot necessarily be trusted, and is reportedly fueling the spread of mis- and disinformation [^1] https://www.axios.com/2023/02/21/chatbots-misinformation-nightmare-chatgpt-ai. Therefore it is imperative to disclose the use of AI in the generation/manipulation of media content. According to the World Economic Forum’s 2024 Global Risks Report (2024, 2024), “misinformation and disinformation” were ranked as the highest global risk anticipated over the next two years. This also comes at a time where OpenAI issues a statement regarding the upcoming 2024 worldwide elections (OpenAI, 2024), emphasizing: “Transparency around AI-generated content”. In fact, in the European Union (EU), such an obligation will be incorporated into the upcoming AI Act (Act, 2022). Despite that the AI Act is still in development and subject to change, the core obligation to “disclose AI-driven interaction” appears to be stable.
@@
Algorithmic approaches to automatically detect text (co-)produced using GenAI tools (e.g., ChatGPT) are so far unreliable (Sadasivan et al., 2023), and can result in consistent biases against specific user groups (e.g., non-native English writers (Liang et al., 2023)). While measures to encode provenance cryptographically in images (cf., Coalition for Content Provenance and Authenticity (C2PA, 2024)) or audio (SynthID (DeepMind, 2024)) are being taken, these are not widely implemented. Moreover, it remains unclear how these need to be displayed to users. Despite concerted efforts worldwide, constructing even simple disclosure measures (e.g., watermarks) that actually succeed in reducing public risks is itself a difficult task. Disclosure is defined as “to make known or public” (Webster, 2024), where in the context of AI, disclosures can “contain information about the data collection, data processing, and decision-making practices of a digital product and are voluntarily provided by the product’s vendor (an individual developer or an organization).” (Open Ethics Initiative (Lukianets, 2024)). A fundamental challenge here is in ensuring effective transparency measures, given the rapid pace of AI development and deployment. To mitigate harms and risks, regulatory efforts such as the AI Disclosure Act of 2023 in the United States (Torres, 2023) and the EU AI Act (Act, 2022) (currently under discussion (of the EU, 2024)) aim to tackle this. In this work, we focus on the EU AI Act proposition (Article 52: “Transparency obligations for providers and users of certain AI systems”), that addresses the issue of AI system transparency [^2] Even though the final text of the AI Act must still be agreed upon, the inclusion of a ”disclosure obligation” is (almost) a certitude..
@@ 2.1. EU AI Act’s Article 52 and AI disclosure implications
We focus on the EU AI Act (Act, 2022) (currently under discussion (of the EU, 2024)) proposition (Article 52: “Transparency obligations for providers and users of certain AI systems”), that addresses AI system transparency. It raises two important clauses:
@@
From these clauses, the types of disclosure are unspecified, and terms such as ‘authentic’ are undefined, leaving them open for interpretation. For example, ‘what’, ‘how’, or ‘when’ should such AI disclosures take place, even if automatically detected, remains unclear. While the most recent draft of the AI Act (Monitor, 2024) touches on aspects of this, where Art. 52(3) (second subparagraph) highlights when to disclose, much remains uncharted regarding human oversight processes when exceptions are made. We believe it is important to carefully assess the AI Act’s disclosure implications, before proceeding to Human-AI interactions, wherein explainability (Panigutti et al., 2023) may be an additional challenge. Furthermore, authenticity is itself a multidimensional concept (Kernis and Goldman, 2006), which raises the question of what aspects of inauthentic media need to be disclosed. AI disclosures, be they automatic (algorithmic) or provided by system designers or other stakeholders, may include the following: revealing when content is AI-generated, registering these emerging AI systems with a database, summarizing copyrighted material used in training these systems, publication of risk assessments, or even trust certification labels (Scharowski et al., 2023). Importantly, encounters with AI-generated content can impact the human experience of algorithms, and more broadly the psychology of Human-AI interaction (Sundar, 2020), to which we turn to next.
@@ 3.2. Workshop materials
The workshop setup consisted of six flip charts that were attached to the walls of the workshop area (see Figure 1(b)). Each flip chart was given a heading of one of the 5W1H terms to indicate to participants which type of questions to ask there. Post-its and markers were provided. Article 52 and Article 3 were printed on an A4 landscape paper, and placed near each flip chart for quick reference. Article 3 provides definitions of ‘provider’ and ‘user’ [^3] In the latest draft version of the EU AI Act (leaked online on January 21, 2024) (Monitor, 2024), ‘user’ is now referred to as ‘deployer’. However, this does not impact our workshop outcomes, even if participants interpreted ‘user’ as intended by the law (‘deployer‘) or as ‘recipient’., that helped participants have a clear understanding of what these ‘legal’ terms mean. Article 3 (current revision; bold added for emphasis) states:
@@ 3.3. Study procedure
Sessions began with a short  10 minute introduction to set the workshop context, to obtain participant consent (data protection and privacy), and for them to fill in a demographics form. Each workshop session consisted of three tasks, each lasting around 25-30 min., amounting to approximately one and a half hours in total. In the introduction, we first presented the EU AI Act, along with Article 52 and a brief explanation of disclosures. Participants were assigned an ID to note down with each post-it note. Participants were split into groups of two or three, depending on the total size in the session. The grouping was only for discussion and logistical purposes, so each question posed was still done individually. Task 1 and 2 were question generation tasks, where participants were instructed to individually generate questions for each of the 5W1H questions. After 10 minutes, they moved to a different question, such that at the end of  60 minutes each group has had the chance to provide their input for all six question types. Participants were also instructed to review the questions that were placed by participants on the flip chart, and to optionally vote with a +/- 1 post-it on the question if they were in agreement/disagreement. For Task 3, we divided each flip chart into two parts: users and providers. This provided participants the opportunity to reflect on and consider whether the posted questions were relevant for the user, the provider, or both (in which case it was placed in the center of the flip chart). Participants were offered candy in the session. When the session ended, participants were thanked for their time and efforts.
@@ 4.4. Theme 4: Trust, Authenticity, and User Empowerment
A key theme (N=33) of our workshops was on trust, authenticity, and how users can be empowered. This consisted of four sub-themes. The first sub-theme includes questions on Authenticity, Provenance, and Transparency: “How important is it for users to know the authenticity of media?[Q81]; “How are we going to verify authenticity?[Q82]; “Where can people manually verify if it is AI generated?[Q84]. The second sub-theme tackles the Human-centered AI issue of Building Trust in AI Systems: “How will different users interpret signals differently?[Q89]; “How does disclosure impact people’s trust and perceived credibility?[Q91]; “How to responsibly disclose artistic uses of AI?[Q92]. The third sub-theme focuses on Empowering Users through Education & Awareness: “How can we train users to spot authentic content without relying on explicit cues?[Q99]; “What type of information empowers/triggers “act of resistance”?[Q100]; “Where should we get ask if we suspect things (images, texts) are generated by AI?[Q103]. Finally, the last sub-theme focuses on User Responsibility and Agency: “Why do users need to know they are interacting with AI?[Q110]; “What information should the users acquire? Are they responsible for it?[Q111]; “Why is it important for promoting democracy?[Q112]. The sub-themes here touch upon important prior work that aims to address authenticity and trust, in the context of media. Liao & Sundar (Liao and Sundar, 2022) introduced a conceptual model called MATCH that describes how trustworthiness is communicated in AI systems through trustworthiness cues, and propose a requirement checklist to support technology creators in identifying reliable cues. Scharowski et al. (Scharowski et al., 2023) examined the potential for AI certification labels (e.g., the ”Digital Trust Label” by the 2023 Swiss Digital Initiative), and found that these can mitigate data-related concerns expressed by end-users (e.g., privacy and data protection), however other concerns (e.g., model performance) remain challenging to address. Longoni et al. (Longoni et al., 2022) found that people rated news headlines written by AI as less accurate than those written by humans. Recently, Toff & Simon (Toff and Simon, 2023) found that on average audiences perceive news labeled as AI-generated as less trustworthy, even when the articles themselves are not evaluated as any less accurate or unfair. However, these effects largely stem from those whose pre-existing levels of trust in news are higher to begin with and among those knowledgeable about journalism.
@@ 5.3. Conclusion
Our work contributes 149 questions clustered into five themes and 18 sub-themes, that we believe can assist in tackling the challenge of transparently communicating AI disclosures – for media providers, as well as end-users consuming media content everyday across platforms and devices. We hope our questions underscore the importance of better understanding user needs and reactions to transparency obligations, and support establishing user-centric designs for transparent AI disclosures that ultimately foster democratic societies based on truth rather than AI-generated fiction.